# PlaceFM: A Training-free Geospatial Foundation Model of Places using Large-Scale Point of Interest Data

### v2

@@ 5.2.1. Population Density Prediction
In this experiment, we leverage the region representations generated by PlaceFM to estimate population density at the ZIP-code level, which serves as a common unit of comparison with baseline methods. Although ZIP codes are chosen here for consistency, our framework is flexible and can be applied to regions of arbitrary granularity. Following baselines, to perform prediction, we train a Random Forest regressor (RF) with 100 decision trees, using 80% of the ZIP-code embeddings for training and the remaining 20% for testing. The ZIP-code level population density data used in our experiments was obtained from publicly available U.S. Census Bureau datasets.
To systematically assess scalability, we evaluate across states of varying sizes. For small-scale settings, we consider Wyoming (WY; 40,165 POIs across 166 ZIP codes) and Vermont (VT; 52,151 POIs across 250 ZIP codes). For medium-scale evaluation, we include Alabama (AL; 255,132 POIs across 626 ZIP codes) and Georgia (GA; 652,234 POIs across 719 ZIP codes). Finally, for large-scale experiments, we assess three diverse and dense states: New York (NY; 1,241,203 POIs across 1,738 ZIP codes), Florida (FL; 1,358,700 POIs across 973 ZIP codes), and California (CA; 2,204,300 POIs across 1,730 ZIP codes). This state selection captures a broad spectrum of graph sizes and densities, enabling a robust demonstration of PlaceFM’s scalability. The complete state-level results are available in the project repository.
Each experiment is repeated 10 times with randomized train-test splits, and average performance is reported. For evaluation, we adopt standard regression metrics: root mean squared error (RMSE), mean absolute error (MAE), and the coefficient of determination $(R^{2})$. Results are summarized in Table 3. For fairness, all baseline models are carefully reimplemented and tuned using their optimal hyperparameter configurations; these settings are reported in Table 4 for reproducibility.
Table 4. Optimal hyperparameter settings for each dataset. For each $\alpha_{i}$, the value shown to the left of the // corresponds to the Population Density prediction task, while the value on the right corresponds to the Housing Price prediction task.
Wyoming (WY) 0.1 Kmeans 0.50.5/0.5 $0.5/0.0$ $0.0/0,0$
Vermont (VT) 0.1 Kmeans $0.0/0.0$ $0.5/0.25$ $0.0/0.0$
Alabama (AL) 0.1 Kmeans $0.0/0.0$ $0.0/0.25$ $0.5/0.0$
Georgia (GA) 0.05 Kmeans $0.5/0.5$ $1.0/1.0$ $0.0/0.0$
New York (NY) 0.02 Kmeans $0.25/0.75$ $0.0/0.0$ $0.0/0.0$
Florida (FL) 0.05 Kmeans $0.0/0.0$ $0.5/0.5$ $1.0/0.0$
California (CA) 0.05 Kmeans $0.0/1.0$ $0.5/0.5$ $1.0/1.0$
Empirical results demonstrate that PlaceFM consistently outperforms baseline approaches across nearly all datasets, underscoring the effectiveness of our training-free region representation learning paradigm. The only exception arises in New York, where PlaceFM ranks second by a marginal gap. We attribute this to the extreme density of the NY graph: when constructing the graph using the region-adaptive Delaunay Triangulation (DT) method, the resulting cluster representations tend to become homogenized, reducing discriminative capacity after POI encoding.
5.2.2. Housing Price Prediction
@@
Housing prices are a critical indicator of both social well-being and economic vitality, and their prediction has long been a central theme in urban studies, economics, and policy research. Accurate housing price estimation not only informs urban planning and economic forecasting but also guides individual decision-making in residential choices. Traditionally, two categories of factors are recognized as most influential: (i) structural attributes of properties, such as size, age, and physical condition, and (ii) locational amenities, which encompass access to services, infrastructure, and neighborhood characteristics. Within the latter category, POIs provide a particularly rich source of information, as they capture the functional and semantic attributes of urban spaces that directly influence housing demand and valuation (Huang et al., 2023).
@@
In this experiment, we leverage the region-level embeddings generated by PlaceFM to predict housing prices at the ZIP-code level, thereby ensuring a consistent basis of comparison with baseline models. Following prior works, we adopt a Random Forest regressor (RF) with 100 decision trees, training on 80% of the ZIP-code embeddings while reserving the remaining 20% for testing. The housing price dataset is obtained from the Zillow Research data repository [^5] https://www.zillow.com/research/data/, specifically the Zillow Home Value Index (ZHVI) for August 2024, which provides a smoothed and standardized measure of typical home values across regions. To comprehensively assess scalability and robustness, we employ the same set of states as in the population density prediction task, encompassing graphs of varying sizes from small (e.g., Wyoming and Vermont) to large (e.g., New York, Florida, and California). Each experiment is repeated ten times with randomized train-test splits, and we report average performance to mitigate sampling variance. Again, for evaluation, we adopt standard regression metrics: RMSE, MAE, and $R^{2}$.
@@
Experimental results, summarized in Table 5, demonstrate that PlaceFM consistently outperforms baseline methods across small- and medium-scale datasets. This highlights the effectiveness of our training-free approach in producing expressive region representations that generalize well across prediction tasks. A notable exception is observed in New York, where PlaceFM ranks third with only a marginal gap from the top-performing models. We believe that this gap arises for the same reasons as in the population density prediction task: the POI embeddings become overly homogenized during encoding, which diminishes their discriminative power.
@@
Figure 4 illustrates the spatial distribution of absolute errors for Vermont and Georgia. The results show that predictions tend to be more accurate in city centers across both states, while larger discrepancies appear in peripheral and expansive regions. In Georgia, we also observe notable errors in highly dense urban areas, particularly in Atlanta. We attribute these patterns to several factors: First, larger regions tend to be more heterogeneous, making their internal variation harder to capture; Second, highly dense areas may exhibit complex housing market dynamics and sharp local variations that are difficult to model.
@@
Figure 4. Spatial distribution of absolute housing price estimation errors. The top figure shows ZIP-code regions in Vermont (VT), and the bottom figure shows those in Georgia (GA).
5.3. Efficiency Comparison
Figure 5. Efficiency comparison of region embedding generation.
@@ 5.3. Efficiency Comparison
Figure 5 illustrates the efficiency and scalability advantages of PlaceFM, which primarily arise from its training-free embedding generation mechanism. For fairness, we exclude Averaging method and PDFM (Agarwal et al., 2024) from this comparison: Averaging incurs negligible computational cost as it simply computes the mean embedding for each region, whereas PDFM relies exclusively on pretrained embeddings from its original implementation. For all other methods, hyperparameters were selected to optimize performance on the downstream population density prediction task described in Section 5.2. For HGI (Huang et al., 2023), the number of training epochs was fixed at 100, fewer than those used in the original study.
@@
As summarized in Figure 5, PlaceFM demonstrates consistent performance while efficiently managing computational resources even as graph size increases. For instance, in Wyoming, the embedding generation time for PlaceFM is 2.41 seconds for $r=0.1$ and 1.32 seconds for $r=0.05$, compared to 34.52 seconds for Place2Vec (Yan et al., 2017) and 60.61 seconds for HGI. Similarly, in Alabama, PlaceFM requires 14.61 seconds ($r=0.1$) and 7.81 seconds ($r=0.05$), while Place2Vec and HGI take 72.12 and 271.24 seconds, respectively. In Florida, the differences are even more pronounced, with PlaceFM completing in 94.65 seconds ($r=0.1$) and 48.29 seconds ($r=0.05$), versus 621.17 seconds for Place2Vec and 1293.84 seconds for HGI. These results indicate that PlaceFM achieves over an order of magnitude faster embedding generation while its runtime grows much more slowly with graph size.
5.4. Transferability Comparison
@@
Table 6. Transferability: ZIP-code embeddings from PlaceFM with different regressor architectures. Units are RMSE, indicating the error of population density estimations (people/km2); lower is better. Best results are in bold.
@@
Place2Vec (Yan et al., 2017) 240.76 246.20 242.12 511.21 498.23 488.25 1115.62 892.45 976.27
HGI (Huang et al., 2023) 170.94 182.55 175.98 458.21 423.05 413.01 772.30 762.43 730.67
PDFM (Agarwal et al., 2024) 308.18 292.11 254.21 536.34 500.07 478.68 806.30 714.09 726.89
PlaceFM (ours) 165.23 171.44 169.12 410.16 401.71 415.12 702.56 710.50 724.00
@@
An essential criterion for assessing urban region representations is their ability to support diverse downstream architectures from a data-centric perspective. Unlike PlaceFM, which produces embeddings through a training-free mechanism, most existing approaches rely on backbone GNNs or other types of neural network architectures for embedding generation. This dependency can introduce inductive biases that limit their flexibility when applied to different downstream models.
@@
Table 6 presents results demonstrating that embeddings derived from PlaceFM generalize robustly across three regression architectures: Random Forest (RF), Multi-Layer Perceptron (MLP, with two hidden layers of 32 and 16 neurons), and Extreme Gradient Boosting (XGB). With the exception of a single case in Georgia, where XGB trained on HGI embeddings achieves the best performance, PlaceFM consistently outperforms competing baselines. Reported values reflect the mean performance over ten runs of downstream inference, following the same setup as Section 5.2, with optimal hyperparameters selected for each model.
5.5. Multi-granular Place Identification
Figure 6. Voronoi spatial distribution of identified places in ZIP code 30329, Atlanta, GA. Places with the same color belong to the same cluster. For visualization clarity, only the top 10 identified places with the largest number of POIs are shown in each column. (a) Clustering using only category features. (b) and (c) Clustering using propagated category features.
Unlike existing geospatial foundation models, which primarily focus on urban region representation learning, PlaceFM pursues a second important objective: the automatic identification of places within each urban region. These places are defined as subgraphs consisting of POIs or higher-level clusters of POIs that are both semantically and spatially similar (see Definition 3.1). By learning to identify such places, PlaceFM enables the discovery of functionally coherent areas within an urban region, such as ZIP codes or cities. This is particularly useful for applications such as place recommendation, where users benefit from being directed toward clusters of POIs that match their preferences, and for urban analytics, where planners and policymakers can assess the functional value of different neighborhoods based on their composition of places (Islam et al., 2022; Zhao et al., 2016). For instance, identifying clusters of restaurants, cafés, and entertainment venues can highlight vibrant social districts, while clusters dominated by schools, libraries, and parks can indicate family-friendly neighborhoods.
Our clustering-based approach reduces the large-scale POI graph of each region into a predefined number of places, given by $k_{r}=\left\lfloor n_{r}\times r\right\rfloor$, where $n_{r}$ is the number of POIs in region $r$ and $r$ is the reduction ratio. The parameter $r$ determines the granularity of the identified places. Larger values of $r$ yield more fine-grained clusters, capturing smaller and more specialized neighborhoods (e.g., a cluster of restaurants and shops around a single intersection), whereas smaller values of $r$ lead to more general and aggregated places that encompass a wider variety of POI types (e.g., a commercial district spanning multiple blocks). In this way, $r$ serves as a natural control parameter for the level of granularity at which urban functionality is represented.
Semantic-only-based Places. In column (a), representations of places are obtained by clustering solely on semantic information, using only category-level POI features. This results in discontinuous and fragmented clusters, where each color corresponds to a semantic category of POIs. For instance, the dark blue cluster represents POIs belonging to category level 1 of Dining and Drinking, which typically includes restaurants, cafés, bars, and similar establishments. This approach overlooks the influence of neighboring POIs in the vicinity of each POI, thereby ignoring valuable spatial context when constructing a representation for a place.
Semantic- and Spatial-Based Places. Unlike in column (a), in columns (b) and (c), PlaceFM leverages both semantic and spatial feature aggregation, producing spatially contiguous and semantically coherent clusters that better reflect meaningful urban places. This property is particularly useful for the traceability of how place embeddings are generated: First, each embedding can be directly linked back to the underlying POIs that constitute the place, along with their spatial configuration and semantic similarity. Such traceability ensures that the learned representations are interpretable, allowing us to understand why two regions are considered functionally similar or different. Second, by aggregating POIs into coherent clusters, we obtain a structured view of the urban landscape, where the relationship between individual places and the overall region becomes explicit. This facilitates not only better transparency in the embedding generation process but also a richer understanding of the functional organization of an entire region.
In column (c), the places with red Voronoi color contain mostly POIs categorized under Dining and Drinking, along with some from Business and Professional Services. This indicates areas that function as mixed-use clusters, where restaurants, cafés, and bars are co-located with offices or service-oriented businesses, reflecting the multifunctional character of urban neighborhoods. In real-world settings, such patterns are commonly found in downtown districts or commercial corridors of large cities, where dining establishments are interwoven with professional offices, coworking spaces, and service providers, creating vibrant hubs of both social and economic activity. .
Effect of reduction ratio $r$. As the reduction ratio $r$ increases, more clusters emerge, capturing finer-grained structures, while smaller values of $r$ highlight broader, more general areas. For example, by doubling $r$, the number of identified places also roughly doubles, but the composition of POIs within each place becomes more unique and specialized. In column (b), compared to column (c), identified places appear more discontinuous and fragmented; however, each place highlights semantically and spatially coherent POIs in greater detail. For instance, in column (b), the green Voronoi areas contain POIs predominantly from category level 1 of Community and Government and Business and Professional Services, which may represent local government offices, community centers, libraries, as well as law firms or consulting offices. Such detailed, specialized clusters are less apparent in the more general, aggregated places shown in column (c), demonstrating how varying $r$ allows control over the granularity of identified urban places and their functional specificity.
6. Conclusion and Future Work
We introduced PlaceFM, a geospatial foundation model that generates general-purpose place embeddings by capturing spatial context and neighborhood structure. Our comprehensive experiments demonstrate its effectiveness in producing region-level embeddings at multiple geographic scales, which can be readily applied to a variety of urban downstream tasks. In this study, the only data modality used for POI features was category information. As future work, PlaceFM could be extended to leverage multi-modal data, including mobility information, to capture human movement patterns, enabling the learning of more informative and meaningful region-level and place-level representations.

### v3

@@ 5.2.1 Population Density Prediction
In this experiment, we use ZIP-code level region representations generated by PlaceRep to estimate population density, enabling direct comparison with baseline methods. Following prior work, we train a Random Forest regressor with 100 decision trees using an 80/20 train-test split over ZIP-code embeddings. Population density labels are obtained from publicly available U.S. Census Bureau datasets. Each experiment is repeated 10 times with randomized train-test splits, and the average performance is reported. Results are summarized in Table 1. For fairness, all baseline models are carefully reimplemented and tuned using their optimal hyperparameter configurations; these settings are reported in Appendix C for reproducibility.
Empirical results demonstrate that PlaceRep consistently outperforms baseline approaches across nearly all datasets, underscoring the effectiveness of our training-free region representation learning paradigm. The only exception arises in New York, where PlaceRep ranks second by a marginal gap. We attribute this to the extreme density of the NY graph: when constructing the graph using the region-adaptive Delaunay Triangulation (DT) method, the resulting cluster representations tend to become homogenized, reducing discriminative capacity after POI encoding.
5.2.2 Housing Price Prediction
@@
Housing prices are a critical indicator of both social well-being and economic vitality, and their prediction has long been a central theme in urban studies, economics, and policy research. Accurate housing price estimation not only informs urban planning and economic forecasting but also guides individual decision-making in residential choices.
@@
In this experiment, we use ZIP-code level embeddings generated by PlaceRep to predict housing prices, ensuring consistency with baseline models. Following prior work, we train a Random Forest regressor with 100 decision trees using an 80/20 train-test split over ZIP-code embeddings. Housing price labels are obtained from the Zillow Research data repository [^2] https://www.zillow.com/research/data/, using the Zillow Home Value Index for August 2024, which provides a standardized measure of typical home values. To evaluate scalability and robustness, we use the same set of states as in the population density task. Each experiment is repeated ten times with randomized splits, and the average performance is reported.
@@
Experimental results, summarized in Table 2, demonstrate that PlaceRep consistently outperforms baseline methods across small- and medium-scale datasets. This highlights the effectiveness of our training-free approach in producing expressive region representations that generalize well across prediction tasks. A notable exception is observed in New York, where PlaceRep ranks third with only a marginal gap from the top-performing models. We believe that this gap arises for the same reasons as in the population density prediction task: the POI embeddings become overly homogenized during encoding, which diminishes their discriminative power. Additional analysis and visualizations of spatial error patterns are provided in Appendix D.1.
5.3 Multi-granular Place Identification
Figure 2: Voronoi spatial distribution of identified places in ZIP code 30329, Atlanta, GA. Places with the same color belong to the same cluster. For visualization clarity, only the top 10 identified places with the largest number of POIs are shown in each column. (a) Clustering using only category features. (b) and (c) Clustering using propagated category features.
PlaceRep aims to automatically identify places within each urban region, defined as subgraphs of POIs or higher-level POI clusters that are both spatially and semantically similar. This enables the discovery of functionally coherent areas within regions such as ZIP codes or cities, supporting applications like place recommendation and urban analytics. For example, clusters of restaurants, cafés, and entertainment venues can reveal vibrant social districts, while clusters dominated by schools, libraries, and parks indicate family-friendly neighborhoods, providing valuable insights for users and urban planners (Islam et al., 2022; Zhao et al., 2016).
Our clustering-based approach reduces the large-scale POI graph of each region into a predefined number of places, given by $k_{r}=\left\lfloor n_{r}\times r\right\rfloor$, where $n_{r}$ is the number of POIs in region $r$ and $r$ is the reduction ratio. The parameter $r$ determines the granularity of the identified places. Larger values of $r$ yield more fine-grained clusters, capturing smaller and more specialized neighborhoods, whereas smaller values of $r$ lead to more general and aggregated places that encompass a wider variety of POI types. In this way, $r$ serves as a natural control parameter for the level of granularity at which urban functionality is represented.
Semantic-only-based Places. In column (a), representations of places are obtained by clustering solely on semantic information, using only category-level POI features. This results in discontinuous and fragmented clusters, where each color corresponds to a semantic category of POIs. This approach overlooks the influence of neighboring POIs in the vicinity of each POI, thereby ignoring valuable spatial context when constructing a representation for a place.
Semantic- and Spatial-Based Places. Unlike column (a), columns (b) and (c) leverage both semantic and spatial feature aggregation, producing spatially contiguous and semantically coherent clusters that better reflect meaningful urban places. This property is critical for traceability: First, each place embedding can be directly linked to its underlying POIs, including their spatial configuration and semantic similarity, enabling interpretable representations and explaining functional similarities or differences between regions. Second, aggregating POIs into coherent clusters provides a structured view of the urban landscape, making the relationship between individual places and the overall region explicit and supporting a richer understanding of regional functional organization.
In column (c), places highlighted in red contain mostly POIs from Dining and Drinking, with some from Business and Professional Services, indicating mixed-use clusters where restaurants, cafés, and bars co-locate with offices and service-oriented businesses. Such patterns commonly occur in downtown districts or commercial corridors, where dining venues are interwoven with professional offices, coworking spaces, and service providers.
Effect of reduction ratio $r$. As the reduction ratio $r$ increases, more clusters emerge, capturing finer-grained structures, while smaller values of $r$ highlight broader, more general areas. In column (b), compared to column (c), identified places appear more discontinuous and fragmented; however, each place highlights semantically and spatially coherent POIs in greater detail. For instance, in column (b), the green Voronoi areas contain POIs predominantly from category level 1 of Community and Government and Business and Professional Services, which may represent local government offices, community centers, libraries, as well as law firms or consulting offices. Such detailed, specialized clusters are less apparent in the more general, aggregated places shown in column (c), demonstrating how varying $r$ allows control over the granularity of identified urban places and their functional specificity.
Additional analyses on the efficiency, and transferability of region embeddings are provided in Appendices D.2 and D.3.
6 Conclusion & Future Work
We introduced PlaceRep, a training-free model that generates general-purpose place embeddings by capturing spatial context and neighborhood structure. Our comprehensive experiments demonstrate its effectiveness in producing region-level embeddings at multiple geographic scales, which can be readily applied to a variety of urban downstream tasks. In this study, the only data modality used for POI features was category information. As future work, PlaceRep could be extended to leverage multi-modal data, including mobility information, to capture human movement patterns, enabling the learning of more informative and meaningful region-level and place-level representations.
Appendix A FSQ-19M POI Graph Dataset
This section provides additional details on the FSQ-19M POI dataset, including preprocessing, statistics, and visualizations. Figure 3 visualizes the spatial distribution of POIs using a uniform random sample of one million entries from FSQ-19M, and Table 3 presents representative examples of spatial and semantic attributes. Section 4.2 in the main paper describes how FSQ-19M is transformed into a graph structure.
Table 3: A data entry in the FSQ-19M dataset.
Attribute Value
longitude 40.858911
latitude -124.073245
state CA
locality Arcata
postcode 95521
date_created 2012-02-08
date_closed None
Table 4: General statistics of the FSQ-19M POI graph dataset
19,019,187 44,202,952 48 31,970 60,155
A.1 Preprocessing and Data Enrichment
This section describes the preprocessing and data enrichment steps used to construct the FSQ-19M dataset from raw Foursquare POI records. For each POI $p$, we retain its latitude ($\text{lat}_{p}$), longitude ($\text{lon}_{p}$), state, locality (city), and ZIP code, which together define its spatial location.
Each POI is also associated with temporal metadata, including its creation date and, when applicable, its closure date, enabling analysis of the temporal evolution of urban environments. In addition, each POI is assigned a hierarchical semantic category defined up to six levels, allowing aggregation and analysis at different semantic granularities.
We apply normalization procedures to textual attributes, such as standardizing state names and ZIP code formats, and remove duplicate entries. To ensure spatial consistency, we perform a spatial join with official U.S. Census ZIP code boundary data to validate and, when necessary, correct ZIP code assignments. The resulting dataset covers a wide range of urban functions, including dining venues, commercial locations, and civic infrastructure, supporting diverse spatial and socioeconomic analyses.
Figure 3: Spatial distribution of POIs based on a uniform random sample of one million entries from the FSQ-19M dataset, covering the 48 contiguous U.S. states.
Appendix B Notations
  • $\mathcal{P}$: Set of all POIs, where each POI $p\in\mathcal{P}$ is represented as $(\text{lat}_{p},\text{lon}_{p},A_{p})$, with $A_{p}=[a_{p_{1}},a_{p_{2}},\dots,a_{p_{K}}]$ denoting the list of $K$ attributes.
  • $a_{p_{i}}$: The $i$-th attribute of POI $p$. In this study, we use the hierarchical category attribute. For example, $\text{category}_{p}$ may be Dining and Drinking $\rightarrow$ Middle Eastern Restaurant $\rightarrow$ Persian Cafe.
  • $G=(\mathcal{V},\mathcal{E})$: POI graph, where $\mathcal{V}=\mathcal{P}$ and edges are defined using a proximity function $d(p_{i},p_{j})$.
  • $G_{r}=(\mathcal{V}_{r},\mathcal{E}_{r})$: Geospatial region represented as a POI subgraph.
Appendix C Implementation Details
This section provides implementation details for PlaceRep and all baseline methods, enabling reproducibility and fair comparison.
C.1 Model Configuration and Training
All baseline methods were reimplemented and tuned based on reported configurations and additional validation under our experimental setting. The full implementation of PlaceRep and all baselines is publicly available at https://github.com/mohammadhashemii/PlaceRep. For downstream tasks, each experiment was repeated ten times using randomized train-test splits, and average performance is reported.
All baseline models except PDFM [Agarwal et al., 2024] were trained from scratch. Due to unavailable architectural details for PDFM, we reused released pretrained embeddings and applied dimensionality reduction to match an embedding size of $d=30$.
For PlaceRep, the feature encoder dimensionality was set to $d=30$. The POI encoder was implemented using a single-layer SGC [Wu et al., 2019] with two propagation steps. Propagation coefficients $\alpha_{0},\alpha_{1},\alpha_{2}$ were tuned within $[0.0,1.0]$ using increments of $0.25$. The clustering module used $k$-means with a maximum of 300 iterations and a convergence threshold of $1\times 10^{-8}$. Full hyper-parameter listing is demonstrated in Table 5.
Table 5: Optimal hyperparameter settings for each dataset. For each $\alpha_{i}$, the value shown to the left of the // corresponds to the Population Density prediction task, while the value on the right corresponds to the Housing Price prediction task.
Wyoming (WY) 0.1 Kmeans 0.50.5/0.5 $0.5/0.0$ $0.0/0,0$
Vermont (VT) 0.1 Kmeans $0.0/0.0$ $0.5/0.25$ $0.0/0.0$
Alabama (AL) 0.1 Kmeans $0.0/0.0$ $0.0/0.25$ $0.5/0.0$
Georgia (GA) 0.05 Kmeans $0.5/0.5$ $1.0/1.0$ $0.0/0.0$
New York (NY) 0.02 Kmeans $0.25/0.75$ $0.0/0.0$ $0.0/0.0$
Florida (FL) 0.05 Kmeans $0.0/0.0$ $0.5/0.5$ $1.0/0.0$
California (CA) 0.05 Kmeans $0.0/1.0$ $0.5/0.5$ $1.0/1.0$
Clustering was executed on Intel Xeon Platinum 8260 CPUs at 2.40GHz using NumPy [Harris et al., 2020]. Baseline training was performed on a high-performance cluster with Tesla A100, V100, and K80 GPUs, selected based on dataset size.
C.2 Baselines
We evaluate PlaceRep against a set of representative baselines, spanning simple heuristics to state-of-the-art geospatial urban region representation learning models, all of which produce region-level embeddings within each state.
2. Place2Vec [Yan et al., 2017]: A method that learns POI category embeddings and then derives region embeddings by averaging over the POIs contained in a region. Unlike simple averaging, Place2Vec incorporates spatial co-occurrence statistics during POI embedding learning, yet the final regional representation is still obtained through aggregation by averaging.
3. HGI [Huang et al., 2023]: An unsupervised model that learns region embeddings by jointly modeling categorical semantics of POIs, POI-level and region-level adjacency, and the interaction between POIs and regions. HGI further employs an attention mechanism to weigh the relative importance of individual POIs during region-level aggregation, enabling more context-sensitive representations.
4. PDFM [Agarwal et al., 2024]: A pre-trained foundation model that integrates diverse geospatial signals, including POI data, maps, activity levels, search trends, weather, and air quality, into a heterogeneous graph. Geographic regions such as counties and postal codes are represented as nodes, with edges reflecting spatial proximity. A GNN is then applied to capture spatial dependencies and generate embeddings for these regions.
Appendix D Additional Experimental Results
D.1 Error Distribution Analysis
@@
Figure 4 illustrates the spatial distribution of absolute errors for Vermont and Georgia for the housing price prediction downstream task. The results show that predictions tend to be more accurate in city centers across both states, while larger discrepancies appear in peripheral and expansive regions. In Georgia, we also observe notable errors in highly dense urban areas, particularly in Atlanta. We attribute these patterns to several factors: First, larger regions tend to be more heterogeneous, making their internal variation harder to capture; Second, highly dense areas may exhibit complex housing market dynamics and sharp local variations that are difficult to model.
@@
Figure 4: Spatial distribution of absolute housing price estimation errors for Vermont (top) and Georgia (bottom).
D.2 Efficiency Comparison
@@
Figure 5 illustrates the efficiency and scalability advantages of PlaceRep, which primarily arise from its training-free embedding generation mechanism. For fairness, we exclude Averaging method and PDFM [Agarwal et al., 2024] from this comparison: Averaging incurs negligible computational cost as it simply computes the mean embedding for each region, whereas PDFM relies exclusively on pretrained embeddings from its original implementation. For all other methods, hyperparameters were selected to optimize performance on the downstream population density prediction task described in Section 5.2. For HGI [Huang et al., 2023], the number of training epochs was fixed at 100, fewer than those used in the original study.
@@
As summarized in Figure 5, PlaceRep demonstrates consistent performance while efficiently managing computational resources even as graph size increases. For instance, in Wyoming, the embedding generation time for PlaceRep is 2.41 seconds for $r=0.1$ and 1.32 seconds for $r=0.05$, compared to 34.52 seconds for Place2Vec [Yan et al., 2017] and 60.61 seconds for HGI. Similarly, in Alabama, PlaceRep requires 14.61 seconds ($r=0.1$) and 7.81 seconds ($r=0.05$), while Place2Vec and HGI take 72.12 and 271.24 seconds, respectively. In Florida, the differences are even more pronounced, with PlaceRep completing in 94.65 seconds ($r=0.1$) and 48.29 seconds ($r=0.05$), versus 621.17 seconds for Place2Vec and 1293.84 seconds for HGI. These results indicate that PlaceRep achieves over an order of magnitude faster embedding generation while its runtime grows much more slowly with graph size.
Figure 5: Efficiency comparison of region embedding generation across methods.
Figure 5 demonstrates that PlaceRep achieves over an order-of-magnitude speedup in embedding generation compared to baseline methods, with runtime scaling more slowly as graph size increases.
D.3 Transferability Comparison
@@
Table 6: Transferability: ZIP-code embeddings from PlaceRep with different regressor architectures. Units are RMSE, indicating the error of population density estimations (people/km2); lower is better. Best results are in bold.
@@
Place2Vec [Yan et al., 2017] 240.76 246.20 242.12 511.21 498.23 488.25 1115.62 892.45 976.27
HGI [Huang et al., 2023] 170.94 182.55 175.98 458.21 423.05 413.01 772.30 762.43 730.67
PDFM [Agarwal et al., 2024] 308.18 292.11 254.21 536.34 500.07 478.68 806.30 714.09 726.89
PlaceRep (ours) 165.23 171.44 169.12 410.16 401.71 415.12 702.56 710.50 724.00
@@
An essential criterion for assessing urban region representations is their ability to support diverse downstream architectures from a data-centric perspective. Unlike PlaceRep, which produces embeddings through a training-free mechanism, most existing approaches rely on backbone GNNs or other types of neural network architectures for embedding generation. This dependency can introduce inductive biases that limit their flexibility when applied to different downstream models.
@@
Table 6 presents results demonstrating that embeddings derived from PlaceRep generalize robustly across three regression architectures: Random Forest (RF), Multi-Layer Perceptron (MLP, with two hidden layers of 32 and 16 neurons), and Extreme Gradient Boosting (XGB). With the exception of a single case in Georgia, where XGB trained on HGI embeddings achieves the best performance, PlaceRep consistently outperforms competing baselines. Reported values reflect the mean performance over ten runs of downstream inference, following the same setup as Section 5.2, with optimal hyperparameters selected for each model.